OpenStudy (anonymous):

Solve the Equation y+5=-12

OpenStudy (anonymous):

@callie2240 @vaibhavm There Ya Go lol

OpenStudy (anonymous):

is it -7

OpenStudy (anonymous):

:) Isolate y so -5 on both sides like i've showed u before :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

iM guessing lol

OpenStudy (anonymous):

pretend the negatives are gone lol 5+12=17 now add the negitives back, -17 would be the answer understand?

OpenStudy (anonymous):

y+5=-12 y=-12-5 Y=17

OpenStudy (anonymous):

-17 do not forget he sign

OpenStudy (anonymous):

Yeah Got It Thanks

OpenStudy (anonymous):

but two - sign is plus (+)

OpenStudy (anonymous):

Haha I Was Way Off

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

only with multiplying @vaibhavm when adding or subtracting its different

OpenStudy (anonymous):

Haha Okay

OpenStudy (anonymous):

lol....anyways we got it......n thanks

OpenStudy (anonymous):

np :) lol

OpenStudy (anonymous):

any more questions...? lol